OpenStudy (anonymous):

Jane has $23 to spend. She buys lunch for $5.47 and a belt for $8.95. Jane then buys a necklace with the remaining money. How much does Jane spend buying the necklace? $3.48 $8.58 $14.05 $14.42

OpenStudy (anonymous):

@priyar

OpenStudy (anonymous):

@misty1212

OpenStudy (cmorris2004):

Hmm another question to easy.

OpenStudy (anonymous):

@CMorris2004 sup bro

OpenStudy (cmorris2004):

Nothing much

OpenStudy (anonymous):

Haha My name is Jane. Moving on. 5.47+8.95= 14.42. 23-14.42= $8.58

OpenStudy (anonymous):

Therefore the answer is B

OpenStudy (anonymous):

thx

OpenStudy (cmorris2004):

i just verified my email.

OpenStudy (cmorris2004):

h

OpenStudy (anonymous):

ok what is it??

OpenStudy (cmorris2004):

thestinger66@gmail.com or you

OpenStudy (anonymous):

xxelfakuxx@gmail.com

OpenStudy (cmorris2004):

youtube gmail type thestinger66

OpenStudy (cmorris2004):

ok illl try that

OpenStudy (anonymous):

i posted my first vid!!

OpenStudy (cmorris2004):

Good

OpenStudy (anonymous):

Which of the following is the solution to the equation x + 6.8 = 53.5? 13.6 33.1 46.7 60.3

OpenStudy (anonymous):

can u like the vid??

OpenStudy (cmorris2004):

Hmm. what your guess

OpenStudy (anonymous):

7.86764705882 is what i got..

OpenStudy (cmorris2004):

Wait here and give me the url for the video

OpenStudy (anonymous):

just go to my channel... in ur Subscriptions

OpenStudy (anonymous):

i invited u to hangouts!!

OpenStudy (anonymous):

Which of the following is the solution to the equation x + 6.8 = 53.5? 13.6 33.1 46.7 60.3

OpenStudy (anonymous):

can u get Skype????

OpenStudy (cmorris2004):

OK and your answer is 46.7. all you do is 53.5-6.8

OpenStudy (cmorris2004):

open your gmail.

OpenStudy (anonymous):

What is the solution to the equation x + 14 = 63? (Input a whole number only.) Numerical Answers Expected! Answer for Blank 1: _________________________________

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

jane wat r u doing here??

OpenStudy (anonymous):

I'm getting messages for replies.

OpenStudy (anonymous):

JANE THE VIRGIN??

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Actually my name is Janie people just call me Jane so....

OpenStudy (anonymous):

What is the solution to the equation x + 14 = 63? (Input a whole number only.) Numerical Answers Expected! Answer for Blank 1: _________________________________

OpenStudy (anonymous):

x=49 =)